OpenStudy (anonymous):

Complex numbers question! Calculate this:

OpenStudy (anonymous):

And represent it on a graph: \[\sqrt[5]{1+i}\]

OpenStudy (anonymous):

go there it may help okay

OpenStudy (anonymous):

https://www.mathway.com/graph

OpenStudy (anonymous):

think of it this way: \(\sqrt[5]{1+i}=a\) then \(a^5=1+i\) or \(a^5=\sqrt2( \cos (\theta+n2\pi) +i sin(\theta +n2\pi))\) where \(\theta=45º \) and n=0,1,2,...\(\infty\) from here, using Moovre's theorem: \(a=\sqrt2(cos\frac{\theta+n2\pi}{5} +i sin\frac{\theta+n2\pi}{5})\) now notice that for n=0,1,2,3,4 you will get different numbers, but startting from 5 and so on they will repeat. So there will be 5 different answers for a. Or to say it oyher way, 5 different roots

OpenStudy (anonymous):

@myko Thank you very much! But can you explain to me once more why we five different root?

OpenStudy (anonymous):

I would guess because a^5 = 1 + i has degree 5...just as every 5th degree polynomial has 5 roots, so would a^5 = 1 + i...

OpenStudy (anonymous):

Ah ok, now I get it. Thank you everyone I really appreciate your help :)

OpenStudy (anonymous):

in the last expretion for \(a\), just substitute n=0, n=1, n=2, n=3, n=4. You will get each time a different expretion. Each of which has argument \(2\pi/5\) bigger tan the previous one. After adding \(2\pi/5 \) five times you will come to same place that you started on the unit sircle. So the numbers will repeat for values higher than 4.

OpenStudy (anonymous):

@naylah